Suppose a,,x,y,z,b are in A.P.when x+y+z=15 and a,l,m,n,b are inH.P.when 1/l+1/m+1/n=5/3.Find a if a>b

Grade:12

1 Answers

Sukhendra Reddy Rompally B.Tech Mining Machinery Engg, ISM Dhanbad
13 years ago
Dear student, since a,x,y,z,b are in AP, a+b = x+z = y = x+y+z/2 = 15/2 ,and since, a,l,m,n,b are in HP,their reciprocals ard in AP, which means, 1/a + 1/b = 1/l + 1/m + 1/n /2 = 5/6 ,that is 6(a+b) = 5ab = 45 ,implies ab=9,a+b=15/2 ,so solve the 2 eqns to get the 2 variables and you get a=6 and b=3/2 THANKS , PLZZ APPROVE MY ANSWER IF YOU LIKE IT